# Mitigating Hallucinations in Vision-Language Models through Image-Guided Head Suppression

### v1

@@ Abstract
Despite their remarkable progress in multimodal understanding tasks, large vision language models (LVLMs) often suffer from “hallucination”, generating texts misaligned with the visual context. Existing methods aimed at reducing hallucinations through inference time intervention incur a significant increase in latency. To mitigate this, we present SPIN, a task-agnostic attention-guided head suppression strategy that can be seamlessly integrated during inference without incurring any significant compute or latency overhead. We investigate whether hallucination in LVLMs can be linked to specific model components. Our analysis suggests that hallucinations can be attributed to a dynamic subset of attention heads in each layer. Leveraging this insight, for each text query token, we selectively suppress attention heads that exhibit low attention to image tokens, keeping the top-K attention heads intact. Extensive evaluations on visual question answering and image description tasks demonstrate the efficacy of SPIN in reducing hallucination scores up to while maintaining F1, and improving throughput by compared to existing alternatives. Code is available here.
@@
Sreetama Sarkar1,††thanks: Equally contributing authors. Yue Che1,∗ Alex Gavin3 Peter A. Beerel1 Souvik Kundu2 1Universiy of Southern California, Los Angeles, USA 2Intel Labs, USA 3Harvard-Westlake High School, Los Angeles, USA {sreetama,yueche,pabeerel}@usc.edu souvikk.kundu@intel.com
@@ 1 Introduction
One key source of hallucination in LVLMs is textual bias, inherited from the pre-trained LLM backbone Liu et al. (2024c); Chen et al. (2024a). To mitigate this, fine-tuning strategies based on Reinforcement Learning (RL) with human or AI-generated feedback have been proposed Sun et al. (2024); Jing and Du (2025). However, these methods are computationally intensive and often impractical for deployment in resource-constrained environments. Recent research has explored inference-only approaches, which modify the decoding pipeline Huang et al. (2024) or apply contrastive decoding techniques Leng et al. (2024); Liu et al. (2024c) to refine logit scores and reduce hallucinations. Although these methods are significantly more efficient compared to training-based approaches, they often struggle to sufficiently reduce hallucinations and suffer from significantly increased latency.
@@
Our Contributions: We perform a detailed analysis revealing that hallucination in LVLMs often stems from specific attention heads exhibiting insufficient attention to visual input. We characterize and quantify these "image-inattentive" heads across model layers. Based on this insight, we propose SPIN, SuPpressing image INattentive heads, a novel, attention-guided head suppression strategy that can be seamlessly integrated during inference, irrespective of the decoding strategy or projection modules. SPIN offers a highly efficient solution since it achieves substantial hallucination reduction without requiring model retraining, and crucially, introduces no additional computational overhead or latency during inference. We evaluate our approach on visual question answering (VQA) and image caption generation tasks. SPIN reduces hallucination scores up to over existing methods, while improving throughput by up to .
@@ 2.2 LVLM Hallucination and Mitigation
Hallucinations can manifest as factual errors to a user’s query or inaccurate image description resulting in non-existent objects, incorrect object attributes or relationships Liu et al. (2024a). Several causes of hallucinations have been identified including biased training data Liu et al. (2023), the inability of vision encoders to accurately ground images Jain et al. (2024), misalignment among different modalities Liu et al. (2024b), and insufficient context attention in LLM decoders Huang et al. (2024); Liu et al. (2024c). Existing hallucination mitigation approaches can be broadly classified into training-based and training-free methods.
@@
Training-based Mitigation: Sun et al. (2024) adapt the Reinforcement Learning from Human Feedback (RLHF), originally developed for text-only models, to the vision-language setting, training VLMs to maximize simulated human rewards. They propose Factually-Augmented RLHF, where the reward model is given access to additional ground truth information, such as image captions, to improve its assessment of factual correctness. In contrast, FGAIF Jing and Du (2025) replaces human supervision with fine-grained AI-generated feedback. It segments model responses into sub-sentences and uses AI models to detect hallucinations relating to object existence, attributes, and relationships. LACING Zhao et al. (2024) addresses language bias in LVLMs by introducing a multimodal dual-attention mechanism and soft-image guidance. It constructs separate attention streams for visual and textual inputs to improve grounding and alignment. Despite their effectiveness, training-based methods are computationally expensive, requiring large-scale resources (e.g., 8× A100 GPUs with 40GB memory each), making them impractical in resource-constrained settings.
@@
Training-free Mitigation: Recently, training-free approaches have been proposed to mitigate the issue of visual context neglect in LVLMs. OPERAHuang et al. (2024) observes that hallucination arises from generating new tokens based on limited summary tokens in which attention aggregation occurs, leading the model to ignore the image context. To address this, OPERA introduces a beam-search variant with a weighted scoring mechanism that downranks candidate sequences exhibiting over-trust patterns. However, its applicability is restricted to beam-search, which is significantly slower than greedy decoding. Other techniques leverage contrastive decoding, which requires multiple forward passes, increasing latency. VCD Leng et al. (2024) builds on the observation that increased visual uncertainty drives models to rely on language priors, thereby amplifying hallucinations. It contrasts the output distributions conditioned on original and distorted visual inputs to identify and suppress hallucinated content. PAI Liu et al. (2024c) mitigates the reduced attention to image tokens through a two-step approach: (1) amplifying attention to image tokens and (2) refining logit scores by subtracting logits computed without image prior to eliminate the text bias. DAMRO Gong et al. (2024) attributes high-attention outlier tokens scattered in the background of the image as the cause for hallucinations and proposes to mitigate the influence of outlier tokens using contrastive decoding. HALC Chen et al. (2024b) is based on contrasting distributions with different visual contexts and using visual matching scores for candidate selection in beam-search. While all the above approaches introduce latency overhead, HALC is reported to cost around 2.4 of the normal greedy decoding time. In contrast, we propose a training-free mitigation strategy that introduces no additional overhead by identifying the role of attention heads in hallucination.
@@ 2.3 Role of Attention Heads
Earlier works Voita et al. (2019) analyzed the role of individual attention heads in neural machine translation, showing that most heads can be pruned without significantly affecting model performance. More recently, routing to expert attention heads Jin et al. (2024); Zheng et al. (2025) was introduced for LLMs and vision transformers, allowing each input token to dynamically select appropriate heads, thereby improving performance. Zhou et al. (2024) further highlighted the impact of specific attention heads on LLM safety, demonstrating that ablating a small subset of heads increases the attack success rate. Recently, Kundu et al. (2025) explored the impact of LLM weight You et al. (2024) and KV quantization Kang et al. (2024) on accuracy as well as hallucination performance of LLMs. However, the role of attention heads in the context of hallucination remains unexplored.
@@ 3 Methodology
Motivational Study: LVLMs struggle with inefficient image attention Chen et al. (2024a); Liu et al. (2024c). In Figure 2, we illustrate how the current token distributes attention across preceding vision and text tokens (constituting system prompt, instruction prompt, and generated output) averaged over all output tokens in the image caption. Our analysis reveals that in deeper layers, generated text tokens allocate 10% of their attention to vision tokens, despite vision tokens comprising 76-92% of the input. This imbalance causes the model to ignore the context provided by the visual input (taken as the “fact”), potentially leading to hallucinations. To mitigate this, for each query text token, we identify a fraction of heads that allocates the least cumulated attention to vision tokens. We then present a training-free strategy to dynamically suppress attention heads, reducing the imbalance in attention and thereby enhancing model performance.
@@
Here, denotes the attention of to input tokens, represents the key matrix for tokens, and denotes the attention score for to the vision tokens only. and denotes the start and end indices of vision tokens. The mask is defined as:
@@
Head and Suppression Factor Selection: To identify the dynamic subset of problematic attention heads for a specific model and task, we adopt an efficient systematic three-stage approach. First, we vary to find the value that achieves the best reduction in hallucinations without a significant drop (3%) in F1. Here, we prune attention heads across all layers equally setting . Adjusting these parameters is explored next. Second, our analysis (Section 4.7) reveals that attention heads in earlier layers tend to contribute more to hallucinations than those in later layers. For specific models and tasks, we find the optimal number of early layers in which we should prune attention heads to minimize hallucination scores. Finally, we explore increasing for the selected heads to mitigate the drop in F1 observed in step 1. In particular, a higher results in improving F1 score at the cost of increasing hallucinations. We choose an that provides an optimal trade-off between hallucinations and F1.
@@ 4.1 Experimental Setup
We evaluate our approach on LLaVA-1.5 (7B, 13B) Liu et al. (2024b), Minigpt4 Zhu et al. (2024), Shikra Chen et al. (2023) and Qwen-VL Bai et al. (2023) models using POPE Li et al. (2023), CHAIR Rohrbach et al. (2018), MMHal-Bench Sun et al. (2024) and GPT-4o assisted evaluation. While POPE and CHAIR evaluate object hallucinations, with POPE using Yes-No questions and CHAIR assessing image descriptions, MMHal-Bench takes into account logical considerations like object count, attributes, and relationships. We report results using different decoding strategies including greedy, nucleus sampling, and beam-search (with beam width of 5), and compare our approach with existing methods like OPERA, VCD, PAI and DAMRO for each decoding strategy. Notably, OPERA is only compatible with beam-search, while VCD is specifically designed for nucleus sampling. We report DAMRO results only for LLaVA models, as its applicability to more complex projection modules like Q-Former remains unclear, a limitation also acknowledged in the original paper. The hyperparameters used for the existing methods for each model are taken from their respective papers (further details are provided in the Appendix A.1). Additionally, to evaluate the efficiency of our method, we report throughput measured in tokens per second. All experiments are run in Pytorch using NVIDIA RTX A6000 GPUs.
@@ 4.2 CHAIR Evaluation
Model | Method | Layers | CS | CI | F1 | ||
LLaVA- 1.5(7B) | Baseline | - | - | - | 44.4 | 12.3 | 77.8 |
PAI | - | - | - | 27.2 | 7.7 | 76.8 | |
DAMRO | - | - | - | 45.2 | 12.4 | 77.8 | |
SPIN | 132 | 0.05 | 0.08 | 26.4 | 7.6 | 77.6 | |
SPIN | 132 | 0.05 | 0.01 | 16.6 | 5.6 | 74.6 | |
LLaVA- 1.5(13B) | Baseline | - | - | - | 41.4 | 10.9 | 78.9 |
PAI | - | - | - | 37.4 | 9.2 | 79.2 | |
DAMRO | - | - | - | 41.2 | 11.0 | 78.7 | |
SPIN | 116 | 0.10 | 0.0 | 30.6 | 8.3 | 79.6 | |
SPIN | 120 | 0.10 | 0.0 | 29.2 | 7.9 | 79.1 | |
Minigpt4 | Baseline | - | - | - | 31.4 | 11.1 | 70.6 |
PAI | - | - | - | 19.8 | 8.4 | 69.7 | |
SPIN | 116 | 0.18 | 0.0 | 21.0 | 6.2 | 68.8 | |
SPIN | 116 | 0.18 | 0.05 | 17.6 | 8.4 | 68.4 | |
Qwen- VL | Baseline | - | - | - | 46.2 | 11.7 | 76.5 |
PAI | - | - | - | 45.0 | 11.3 | 76.6 | |
SPIN | 120 | 0.30 | 0.08 | 29.6 | 8.5 | 78.0 | |
SPIN | 120 | 0.30 | 0.001 | 24.8 | 7.8 | 76.7 | |
Shikra | Baseline | - | - | - | 55.2 | 14.0 | 75.39 |
PAI | - | - | - | 34.0 | 8.5 | 76.4 | |
SPIN | 132 | 0.40 | 0.0 | 30.6 | 8.0 | 75.41 | |
SPIN | 132 | 0.45 | 0.0 | 24.4 | 7.1 | 74.3 |
Model | Mode | Layers | Supp. | Scale | Random | Popular | Adversarial | Overall | ||||
Heads | Factor | Accuracy | F1 | Accuracy | F1 | Accuracy | F1 | Accuracy | F1 | |||
LLaVA-1.5 (7B) | Baseline | - | - | - | 86.77 | 85.21 | 85.73 | 84.36 | 84.57 | 83.27 | 85.69 | 84.28 |
PAI | - | - | - | 88.53 | 87.54 | 87.40 | 86.53 | 85.43 | 84.59 | 87.12 | 86.22 | |
DAMRO | - | - | - | 86.73 | 85.18 | 85.77 | 84.39 | 84.53 | 83.22 | 85.68 | 84.26 | |
SPIN | 132 | 0.20 | 0.1 | 89.47 | 88.66 | 88.53 | 87.84 | 86.63 | 85.89 | 88.21 | 87.46 | |
LLaVA-1.5 (13B) | Baseline | - | - | - | 86.77 | 85.81 | 86.27 | 85.56 | 82.33 | 82.07 | 85.12 | 84.48 |
PAI | - | - | - | 87.67 | 87.03 | 86.97 | 86.60 | 83.03 | 83.14 | 85.89 | 85.59 | |
SPIN | 120 | 0.35 | 0.0 | 91.17 | 90.65 | 88.83 | 88.46 | 85.53 | 85.49 | 88.51 | 88.20 | |
Minigpt4 | Baseline | - | - | - | 83.35 | 81.28 | 75.23 | 73.79 | 75.66 | 73.52 | 78.08 | 76.20 |
PAI | - | - | - | 85.25 | 83.62 | 75.97 | 74.93 | 77.91 | 76.10 | 79.71 | 78.22 | |
SPIN | 124 | 0.05 | 0.0 | 85.64 | 84.56 | 75.36 | 75.34 | 77.50 | 76.86 | 79.50 | 78.92 | |
Shikra | Baseline | - | - | - | 80.77 | 80.73 | 78.06 | 78.71 | 75.99 | 77.22 | 78.27 | 78.89 |
PAI | - | - | - | 80.73 | 80.30 | 77.90 | 78.52 | 75.90 | 77.08 | 78.18 | 78.63 | |
SPIN | 116 | 0.20 | 0.001 | 82.04 | 81.56 | 79.15 | 79.38 | 76.78 | 77.56 | 79.32 | 79.50 |
@@
CHAIR compiles a set of ground-truth objects for each image. Any object included in the image caption that is not present in the ground-truth object set is classified as a “hallucinated object”. We evaluate our approach on randomly sampled 500 images from the COCO 2014 Lin et al. (2014) validation set with the prompt “Please help me describe the image in detail.”, following the same setup as used in Liu et al. (2024c); Huang et al. (2024). We report F1 scores alongside CHAIR scores to ensure that reducing hallucinations does not come at the expense of missing correct objects.
@@
Evaluation on greedy decoding: We present CHAIR results for Baseline, PAI, DAMRO and SPIN using Greedy decoding in Table 1. A lower CHAIR score indicates fewer hallucinated objects and a higher F1 score indicates higher caption accuracy. SPIN reduces Ci and Cs by 2.2 and 2.7 respectively over baseline LLaVA 7B with 3% degradation in F1 when 5% heads are suppressed uniformly across all layers. The drop in F1 can be mitigated using a higher scaling factor while still maintaining lower CHAIR scores compared to existing methods. For LLaVA 13B, Qwen-VL and Shikra, SPIN outperforms baseline even in terms of F1 scores while reducing Ci and Cs by 1.38 and 1.42 for LLaVA 13B, 1.5 and 1.9 for Qwen-VL, and 1.97 and 2.3 for Shikra. For Minigpt4, SPIN reduces Ci and Cs by 1.8 and 1.5 with 1.8% degradation in F1. We also outperform PAI and DAMRO for all models using greedy decoding. These significant improvements in the CHAIR scores quantify the benefits of the head pruning.
@@
In general, we observe that the problematic heads lie in the first 16 to 20 layers, whereas for some models like LLaVA 7B and Shikra, they are uniformly distributed across all layers. The fraction of suppressed heads () varies significantly across models, ranging from 5% in LLaVA-7B to 45% in Shikra. Shikra’s higher head suppression ratio likely indicates its attention heads perform more distributed or redundant functions, as opposed to LLaVa, where heads are more specialized. A detailed ablation study analyzing the impact of layer position, the fraction of suppressed heads, and the choice of scaling factor is provided in Section 4.7.
@@
Evaluation on other decoding strategies: In Figure 3, we compare SPIN to existing methods for different decoding modes for LLaVA 7B and Qwen-VL. We compare with OPERA for beam-search decoding, VCD for nucleus sampling, DAMRO for greedy and nucleus sampling, and PAI for all three modes. Our results demonstrate that we achieve up to 3 lower CHAIR scores compared to existing methods for greedy and beam-search. For nucleus sampling, SPIN exhibits a slight degradation in performance likely due to the variability in token selection, as further elaborated in Section 4.

### v2

@@ Abstract
Despite their remarkable progress in multimodal understanding tasks, large vision language models (LVLMs) often suffer from “hallucination”, generating texts misaligned with the visual context. Existing methods aimed at reducing hallucinations through inference time intervention incur a significant increase in latency. To mitigate this, we present SPIN, a task-agnostic attention-guided head suppression strategy that can be seamlessly integrated during inference without incurring any significant compute or latency overhead. We investigate whether hallucination in LVLMs can be linked to specific model components. Our analysis suggests that hallucinations can be attributed to a dynamic subset of attention heads in each layer. Leveraging this insight, for each text query token, we selectively suppress attention heads that exhibit low attention to image tokens, keeping the top- attention heads intact. Extensive evaluations on visual question answering and image description tasks demonstrate the efficacy of SPIN in reducing hallucination scores up to
while maintaining F1, and improving throughput by
compared to existing alternatives. Code is available here.
@@
Sreetama Sarkar1,††thanks: Equally contributing authors. Yue Che1,∗ Alex Gavin3 Peter A. Beerel1 Souvik Kundu21University of Southern California, Los Angeles, USA 2Intel Labs, USA3Harvard-Westlake School, Los Angeles, USA{sreetama,yueche,pabeerel}@usc.edu souvikk.kundu@intel.com
@@ 1 Introduction
Figure 1: Caption generation using LLaVA-1.5 and SPIN. LLaVA-1.5’s generated text description mentions a “chair” in the background, which is clearly a hallucinated object. SPIN mitigates hallucination while successfully identifying the objects present in the image.
One key source of hallucination in LVLMs is textual bias, inherited from the pre-trained LLM backbone Liu et al. (2024d); Chen et al. (2024a). To mitigate this, fine-tuning strategies based on Reinforcement Learning (RL) with human or AI-generated feedback have been proposed Sun et al. (2024); Jing and Du (2025). However, these methods are computationally intensive and often impractical for deployment in resource-constrained environments. Recent research has explored inference-only approaches, which modify the decoding pipeline Huang et al. (2024) or apply contrastive decoding techniques Leng et al. (2024); Liu et al. (2024d) to refine logit scores and reduce hallucinations. Although these methods are significantly more efficient compared to training-based approaches, they often struggle to sufficiently reduce hallucinations and suffer from significantly increased latency.
@@
Our Contributions:
We perform a detailed analysis revealing that hallucination in LVLMs often stems from specific attention heads exhibiting insufficient attention to visual input. We characterize and quantify these "image-inattentive" heads across model layers. Based on this insight, we propose SPIN, SuPpressing image INattentive heads, a novel, attention-guided head suppression strategy that can be seamlessly integrated during inference, irrespective of the decoding strategy or projection modules. SPIN offers a highly efficient solution since it achieves substantial hallucination reduction without requiring model retraining, and crucially, introduces no additional computational overhead or latency during inference.
We evaluate our approach on visual question answering (VQA) and image caption generation tasks. SPIN reduces hallucination scores up to over existing methods, while improving throughput by up to .
@@ 2.2 LVLM Hallucination and Mitigation
Hallucinations can manifest as factual errors to a user’s query or inaccurate image description resulting in non-existent objects, incorrect object attributes or relationships Liu et al. (2024a).
Several causes of hallucinations have been identified including biased training data Liu et al. (2023), the inability of vision encoders
to accurately ground images Jain et al. (2024), misalignment among different
modalities Liu et al. (2024b), and insufficient context attention in LLM decoders Huang et al. (2024); Liu et al. (2024d).
Existing hallucination mitigation approaches can be broadly classified into training-based and training-free methods.
Training-based Mitigation:
Sun et al. (2024) adapt the Reinforcement
@@
Learning from Human Feedback (RLHF), originally developed for text-only models, to the vision-language setting, training VLMs to maximize simulated human rewards. They propose Factually-Augmented RLHF, where the reward model is given access to additional ground truth information, such as image captions, to improve its assessment of factual correctness. In contrast, FGAIF Jing and Du (2025) replaces human supervision with fine-grained AI-generated feedback. It segments model responses into sub-sentences and uses AI models to detect hallucinations relating to object existence, attributes, and relationships.
LACING Zhao et al. (2024) addresses language bias in LVLMs by introducing a multimodal dual-attention mechanism and soft-image guidance. It constructs separate attention streams for visual and textual inputs to improve grounding and alignment. Despite their effectiveness, training-based methods are computationally expensive, requiring large-scale resources (e.g., 8× A100 GPUs with 40GB memory each), making them impractical in resource-constrained settings.
Training-free Mitigation:
Recently, training-free approaches
@@
have been proposed to mitigate the issue of visual context neglect in LVLMs. OPERAHuang et al. (2024) observes that hallucination arises from generating new tokens based on limited summary tokens in which attention aggregation occurs, leading the model to ignore the image context. To address this, OPERA introduces a beam-search variant with a weighted scoring mechanism that downranks candidate sequences exhibiting over-trust patterns. However, its applicability is restricted to beam-search, which is significantly slower than greedy decoding. Other techniques leverage contrastive decoding, which requires multiple forward passes, increasing latency.
VCD Leng et al. (2024) builds on the observation that increased visual uncertainty drives models to rely on language priors, thereby amplifying hallucinations. It contrasts the output distributions conditioned on original and distorted visual inputs to identify and suppress hallucinated content.
PAI Liu et al. (2024d) mitigates the reduced attention to image tokens through a two-step approach: (1) amplifying attention to image tokens and (2) refining logit scores by subtracting logits computed without image prior to eliminate the text bias.
DAMRO Gong et al. (2024) attributes high-attention outlier tokens scattered in the background of the image as the cause for hallucinations and proposes to mitigate the influence of outlier tokens using contrastive decoding.
HALC Chen et al. (2024b) is based on contrasting distributions with different visual contexts and using visual matching scores for candidate selection in beam-search. While all the above approaches introduce latency overhead, HALC is reported to cost around 2.4 of the normal greedy decoding time. Recently, Liu et al. (2025) introduced a test-time latent space steering approach through visual and textual intervention (VTI) using pre-computed steering vectors, whereas ICT Chen et al. (2025) proposes image and object-level intervention to apply targeted activation shifts to selected attention heads identified through binary classifiers, requiring additional parameters.
In contrast, we propose a training-free mitigation strategy that introduces no additional overhead by identifying the role of attention heads in hallucination.
@@ 2.3 Role of Attention Heads
Earlier works
Voita et al. (2019) analyzed the role of individual attention heads in neural machine translation, showing that most heads can be pruned without significantly affecting model performance. More recently, routing to expert attention heads Jin et al. (2024); Zheng et al. (2025) was introduced for LLMs and vision transformers, allowing each input token to dynamically select appropriate heads, thereby improving performance. Zhou et al. (2024) further highlighted the impact of specific attention heads on LLM safety, demonstrating that ablating a small subset of heads increases the attack success rate. Recently, Kundu et al. (2025) explored the impact of LLM weight you (2024) and KV quantization Kang et al. (2024) on accuracy as well as hallucination performance of LLMs. However, the role of attention heads in the context of hallucination remains unexplored.
Figure 2: Average attention allocated by the current token to the preceding vision and text tokens in CHAIR. Image tokens receive 10% of total attention from layer 3, while constituting 76-92% of the input.
@@ 3 Methodology
Motivational Study:
LVLMs struggle with inefficient image attention Chen et al. (2024a); Liu et al. (2024d). In Figure 2, we illustrate how the current token distributes attention across preceding vision and text tokens (constituting system prompt, instruction prompt, and generated output) averaged over all output tokens in the image caption. Our analysis reveals that in deeper layers, generated text tokens allocate 10% of their attention to vision tokens, despite vision tokens comprising 76-92% of the input. This imbalance causes the model to ignore the context provided by the visual input (taken as the “fact”), potentially leading to hallucinations. To mitigate this, for each query text token, we identify a fraction of heads that allocates the least cumulated attention to vision tokens.
We then present a training-free strategy to dynamically suppress attention heads, reducing the imbalance in attention and thereby enhancing model performance.
@@
Here, denotes the attention of to input tokens, represents the key matrix for tokens, and denotes the attention score for to the vision tokens only. and denotes the start and end indices of vision tokens.
The mask is defined as:
@@
Head and Suppression Factor Selection: To identify the dynamic subset of problematic attention heads for a specific model and task, we adopt an efficient systematic three-stage approach. First, we vary to find the value that achieves the best reduction in hallucinations without a significant drop (3%) in F1. Here, we prune attention heads across all layers equally setting . Adjusting these parameters is explored next.
Second, our analysis (Section 4.8) reveals that attention heads in earlier layers tend to contribute more to hallucinations than those in later layers. For specific models and tasks, we find the optimal number of early layers in which we should prune attention heads to minimize hallucination scores.
Finally, we explore increasing for the selected heads to mitigate the drop in F1 observed in step 1. In particular, a higher results in improving F1 score at the cost of increasing hallucinations. We choose an that provides an optimal trade-off between hallucinations and F1.
@@ 4 Experimental Results
Figure 3: CHAIR scores for SPIN compared with existing approaches for greedy, beam-search, and nucleus sampling based decoding on LLaVA-1.5 (7B) (top) and Qwen-VL (bottom).
@@ 4.1 Experimental Setup
We evaluate our approach on LLaVA-1.5 (7B, 13B) Liu et al. (2024b), LLaVA-Next Liu et al. (2024c), MiniGPT-4 Zhu et al. (2024), Shikra Chen et al. (2023), Qwen-VL Bai et al. (2023), and Qwen2.5-VL Bai et al. (2025) models using POPE Li et al. (2023), CHAIR Rohrbach et al. (2018), MMHal-Bench Sun et al. (2024), MME Fu et al. (2023), MMMU Yue et al. (2024), and GPT-4o assisted evaluation. While POPE and CHAIR evaluate object hallucinations, with POPE using Yes-No questions and CHAIR assessing image descriptions, MMHal-Bench takes into account logical considerations like object count, attributes, and relationships. We report results using different decoding strategies including greedy, nucleus sampling, and beam-search (with beam width of 5), and compare our approach with existing methods like OPERA, VCD, PAI, and DAMRO for each decoding strategy. Notably, OPERA is only compatible with beam-search, while VCD is specifically designed for nucleus sampling. We report DAMRO results only for LLaVA models, as its applicability to more complex projection modules like Q-Former remains unclear, a limitation also acknowledged in the original paper. The hyperparameters used for the existing methods for each model are taken from their respective papers (further details are provided in the Appendix A.1). Additionally, to evaluate the efficiency of our method, we report throughput measured in tokens per second.
All experiments are run in Pytorch using NVIDIA RTX A6000 GPUs.
@@ 4.2 CHAIR Evaluation
Model
Method
Layers
CS
CI
F1
LLaVA-
1.5 (7B)
Baseline
44.4
12.3
77.8
PAI
27.2
7.7
76.8
DAMRO
45.2
12.4
77.8
VTI
35.8
11.1
76.8
SPIN
132
0.05
0.08
26.4
7.6
77.6
SPIN
132
0.05
0.01
16.6
5.6
74.6
LLaVA-
1.5 (13B)
Baseline
41.4
10.9
78.9
PAI
37.4
9.2
79.2
DAMRO
41.2
11.0
78.7
SPIN
116
0.10
0.0
30.6
8.3
79.6
SPIN
120
0.10
0.0
29.2
7.9
79.1
MiniGPT-4
Baseline
31.4
11.1
70.6
PAI
19.8
8.4
69.7
SPIN
116
0.18
0.0
21.0
6.2
68.8
SPIN
116
0.18
0.05
17.6
8.4
68.4
Qwen-VL
Baseline
46.2
11.7
76.5
PAI
45.0
11.3
76.6
SPIN
120
0.30
0.08
29.6
8.5
78.0
SPIN
120
0.30
0.001
24.8
7.8
76.7
Shikra
Baseline
55.2
14.0
75.39
PAI
34.0
8.5
76.4
SPIN
132
0.40
0.0
30.6
8.0
75.41
SPIN
132
0.45
0.0
24.4
7.1
74.3
Table 1: CHAIR evaluation with Greedy decoding.
Model
Mode
Layers
Supp.
Scale
Random
Popular
Adversarial
Overall
Heads
Factor
Accuracy
F1
Accuracy
F1
Accuracy
F1
Accuracy
F1
LLaVA-1.5
Baseline
86.77
85.21
85.73
84.36
84.57
83.27
85.69
84.28
PAI
88.53
87.54
87.40
86.53
85.43
84.59
87.12
86.22
DAMRO
86.73
85.18
85.77
84.39
84.53
83.22
85.68
84.26
SPIN
132
0.20
0.1
89.47
88.66
88.53
87.84
86.63
85.89
88.21
87.46
LLaVA-1.5
Baseline
86.77
85.81
86.27
85.56
82.33
82.07
85.12
84.48
PAI
87.67
87.03
86.97
86.60
83.03
83.14
85.89
85.59
SPIN
120
0.35
0.0
91.17
90.65
88.83
88.46
85.53
85.49
88.51
88.20
MiniGPT-4
Baseline
83.35
81.28
75.23
73.79
75.66
73.52
78.08
76.20
PAI
85.25
83.62
75.97
74.93
77.91
76.10
79.71
78.22
SPIN
124
0.05
0.0
85.64
84.56
75.36
75.34
77.50
76.86
79.50
78.92
Shikra
Baseline
80.77
80.73
78.06
78.71
75.99
77.22
78.27
78.89
PAI
80.73
80.30
77.90
78.52
75.90
77.08
78.18
78.63
SPIN
116
0.20
0.001
82.04
81.56
79.15
79.38
76.78
77.56
79.32
79.50
Table 2: Multi-turn POPE evaluation across random, popular, and adversarial splits using Greedy Decoding.
@@
CHAIR compiles a set of ground-truth objects for each image. Any object included in the image caption that is not present in the ground-truth object set is classified as a “hallucinated object”. We evaluate our approach on randomly sampled 500 images from the COCO 2014 Lin et al. (2014) validation set with the prompt “Please help me describe the image in detail.”, following the same setup as used in Liu et al. (2024d); Huang et al. (2024). We report F1 scores alongside CHAIR scores to ensure that reducing hallucinations does not come at the expense of missing correct objects.
Evaluation on greedy decoding: We present CHAIR results for Baseline, PAI, DAMRO, VTI, and SPIN using Greedy decoding in Table 1. A lower CHAIR score indicates fewer hallucinated objects and a higher F1 score indicates higher caption accuracy. SPIN reduces Ci and Cs by
2.2 and
@@
2.7 respectively over baseline LLaVA 7B with 3% degradation in F1 when 5% heads are suppressed uniformly across all layers. The drop in F1 can be mitigated using a higher scaling factor while still maintaining lower CHAIR scores compared to existing methods. For LLaVA 13B, Qwen-VL and Shikra, SPIN outperforms baseline even in terms of F1 scores while reducing Ci and Cs by 1.38 and 1.42 for LLaVA 13B, 1.5 and 1.9 for Qwen-VL, and 1.97 and 2.3 for Shikra. For MiniGPT-4, SPIN reduces Ci and Cs by 1.8 and 1.5 with 1.8% degradation in F1. We also outperform PAI, DAMRO, and VTI for all models using greedy decoding. These significant improvements in the CHAIR scores quantify the benefits of the head pruning. We provide additional CHAIR evaluation results for LLaVA-Next and Qwen2.5-VL in the Appendix Table 12.
@@
In general, we observe that the problematic heads lie in the first 16 to 20 layers, whereas for some models like LLaVA 7B and Shikra, they are uniformly distributed across all layers. The fraction of suppressed heads () varies significantly across models, ranging from 5% in LLaVA 7B to 45% in Shikra. Shikra’s higher head suppression ratio likely indicates its attention heads perform more distributed or redundant functions, as opposed to LLaVA, where heads are more specialized. A detailed ablation study analyzing the impact of layer position, the fraction of suppressed heads, and the choice of scaling factor is provided in Section 4.8.
Figure 4: MMHal-Bench evaluation on LLaVA-1.5 (13B), MiniGPT-4 and Shikra.
Evaluation on other decoding strategies:
@@
In Figure 3, we compare SPIN to existing methods for different decoding modes for LLaVA 7B and Qwen-VL. We compare with OPERA for beam-search decoding, VCD for nucleus sampling, DAMRO for greedy and nucleus sampling, and PAI for all three modes. Our results demonstrate that we achieve up to 3 lower CHAIR scores compared to existing methods for greedy and beam-search. For nucleus sampling, SPIN exhibits a slight degradation in performance likely due to the variability in token selection, as further elaborated in Section 4.
Mitigating Repetition from Head Pruning: